# Error Slice Discovery via Manifold Compactness

### v1

@@ Abstract
Despite the great performance of deep learning models in many areas, they still make mistakes and underperform on certain subsets of data, i.e. error slices. Given a trained model, it is important to identify its semantically coherent error slices that are easy to interpret, which is referred to as the error slice discovery problem. However, there is no proper metric of slice coherence without relying on extra information like predefined slice labels. Current evaluation of slice coherence requires access to predefined slices formulated by metadata like attributes or subclasses. Its validity heavily relies on the quality and abundance of metadata, where some possible patterns could be ignored. Besides, current algorithms cannot directly incorporate the constraint of coherence into their optimization objective due to the absence of an explicit coherence metric, which could potentially hinder their effectiveness. In this paper, we propose manifold compactness, a coherence metric without reliance on extra information by incorporating the data geometry property into its design, and experiments on typical datasets empirically validate the rationality of the metric. Then we develop Manifold Compactness based error Slice Discovery (MCSD), a novel algorithm that directly treats risk and coherence as the optimization objective, and is flexible to be applied to models of various tasks. Extensive experiments on the benchmark and case studies on other typical datasets demonstrate the superiority of MCSD.
Machine Learning, ICML
[…]
@@ 1 Introduction
In recent years, with the enhancement of computational power, neural networks have achieved significant progress in numerous tasks (He et al., 2016; Devlin et al., 2018). Despite their impressive overall performance, they are far from perfect, and still suffer from performance degradation on some subpopulations (Yang et al., 2023). This substantially hinders their application in risk-sensitive scenarios like medical imaging (Suzuki, 2017), autonomous driving (Huval et al., 2015), etc., where model mistakes may result in catastrophic consequences. Therefore, to avoid the misuse of models, it is a fundamental problem to identify subsets (or slices) where a given model tends to underperform. Moreover, we would like to find coherent interpretable semantic patterns in the underperforming slices. For example, a facial recognition model may underperform in certain demographic groups like elderly females. An autonomous driving system may fail in the face of steep road conditions. Identifying such coherent patterns could help us understand model failures, and we could employ straightforward solutions for improvement like collecting new data (Liu et al., 2023) or upweighting samples in error slices (Liu et al., 2021a).
@@
Previously, there are works of error slice discovery (d’Eon et al., 2022; Eyuboglu et al., 2022; Wang et al., 2023b) towards this goal. Despite the emphasis on coherence in error slice discovery, there is no proper metric to assess the coherence of a given slice without additional information like predefined slice labels. On one hand, this impairs the efficacy of the evaluation paradigm of error slice discovery. In the current benchmark (Eyuboglu et al., 2022), with the help of metadata like attributes or subclasses, it predefines slices that are already semantically coherent, and they depict the coherence of a slice discovered by a specific algorithm via the matching degrees between it and the predefined underperforming slices, so as to evaluate the effectiveness of the algorithm. Such practice heavily relies on not only the availability but also the quality of metadata, whose annotations are usually expensive, and may overlook model failure patterns not captured by existing metadata. On the other hand, due to the absence of an explicit coherence metric, current algorithms can only indirectly incorporate the constraint of coherence into their design, e.g. via clustering (Eyuboglu et al., 2022; Wang et al., 2023b; Plumb et al., 2023), without treating it as a direct optimization objective. This could potentially impede the development of more effective error slice discovery algorithms.
@@
Figure 1: Illustration of MCSD. The blue points are correctly classified by the given trained model, while the red ones are wrongly classified. We can see that the trained model achieves a good overall accuracy, but exhibit a high error in a certain slice.
@@
In this paper, inspired by the data geometry property that high dimensional data tends to lie on a low-dimensional manifold (Belkin & Niyogi, 2003; Roweis & Saul, 2000; Tenenbaum et al., 2000), we incorporate this property to propose manifold compactness as the metric of coherence given a slice, which does not require additional information. We illustrate the validity of the metric by showing that it captures semantic patterns better than depicting coherence via metrics directly calculated in Euclidean space, and is empirically consistent with current evaluation metrics that require predefined slice labels. Then we propose a novel and flexible algorithm named Manifold Compactness based error Slice Discovery (MCSD) that jointly optimizes the average risk and manifold compactness to identify the error slice. Thus both the risk and coherence, i.e. the desired properties of error slices are explicitly treated as the optimization objective. We illustrate our algorithm in Figure 1. Besides, our algorithm can be directly applied to trained models of different tasks while most error slice discovery methods are restricted to classification only. We provide theoretical analyses of the algorithm. We conduct experiments on dcbench (Eyuboglu et al., 2022) to demonstrate our algorithm’s superiority compared with existing ones. We also provide several case studies on different types of datasets and tasks to showcase the effectiveness and flexibility of our algorithm. Our contributions are summarized below:
@@ 2 Problem
Unless stated otherwise, for random variables, uppercase letters are used, in contrast to a concrete dataset where lowercase letters are used. Consider a general setting of supervised learning. The input variable is denoted as $X\in\mathcal{X}$ and the outcome is denoted as $Y\in\mathcal{Y}$, whose joint distribution is $P(X,Y)$. There exist multiple slices, where $j$-th slice can be represented as a slice label variable $S^{(j)}\in\{0,1\}$. For the classic supervised learning, the goal is to learn a model $f_{\theta}:\mathcal{X}\mapsto\mathcal{Y}$ with parameter $\theta$. Denote $\ell:\mathcal{Y}\times\mathcal{Y}\mapsto[0,+\infty]$ as the loss function. Current machine learning algorithms are capable of learning models with a satisfying overall performance, which can be demonstrated via a low risk $\mathbb{E}_{P}[\ell(f_{\theta}(X),Y)]$ over the whole population. However, performance degradation could still occur in a certain subpopulation or slice. Here we introduce the error slice discovery problem:
@@
The reason why we require an extra validation dataset to implement error slice discovery is that for deep learning models, training data is usually fitted well enough or even nearly perfect. Thus model mistakes on training data carry much less information on models’ generalization ability. This is common practice in previous works (d’Eon et al., 2022; Eyuboglu et al., 2022; Wang et al., 2023b). Without ambiguity, we omit the superscript or subscript of “va” for $n,x_{i},y_{i}$ for convenience in the next two sections.
@@ 3 Metric
Due to the absence of a proper metric for coherence that is independent of additional information, the current benchmark (Eyuboglu et al., 2022) provides numerous datasets, trained models, and their predefined underperforming slice labels. They employ precision@$k$, i.e. the proportion of the top $k$ elements in the discovered slice belonging to the predefined ground-truth error slice as the metric of slice coherence to evaluate error slice discovery algorithms. Although such practice is reasonable to some extent, its effectiveness of evaluation strongly relies on the quality of metadata that composes the underperforming slice labels, which might be not even available under many circumstances.
@@
Figure 2: Category “Blond Hair” of CelebA. Visualization of t-SNE and UMAP (manifold-based dimension reduction techniques) shows much clearer clustering structures than that of PCA (mainly preserving Euclidean distances between data points), indicating that it could be better to measure coherence in the metric space of a manifold than using metrics directly calculated in Euclidean space.
To eliminate the requirement of predefined slices, we try to propose a new metric of coherence. It is commonly acknowledged that high-dimensional data usually lies on a low-dimensional manifold (Belkin & Niyogi, 2003; Roweis & Saul, 2000; Tenenbaum et al., 2000). In this case, while direct usage of Euclidean distance cannot properly capture the dissimilarity between data points, the geodesic distance in the metric space of the manifold can. For preliminary justification, here we provide visualization analyses based on different types of dimension-reduction techniques. Among these techniques, PCA mainly preserves pairwise Euclidean distances between data points while t-SNE and UMAP are both manifold learning techniques. In Figure 2, blue dots are correctly classified by the trained model and red dots are wrongly classified. We can see that the visualization of t-SNE and UMAP shows much clearer clustering structures than that of PCA, either having a larger number of clusters or exhibiting larger margins between clusters. This indicates that it could be better to measure coherence in the metric space of a manifold than in the original Euclidean space. Due to the space limit, we only present results of the widely adopted facial dataset CelebA (Liu et al., 2015) here, leaving results of other datasets in Section A.1.1, where the same conclusion is true.
Therefore, we attempt to define a metric of coherence inside the discovered slice via compactness in the data manifold. In practice, a manifold can be treated as a graph $G$ (Melas-Kyriazi, 2020), and we can apply graph learning methods like k-nearest neighbor (kNN) to approximate it (Dann et al., 2022). Given an identified slice $\hat{\mathcal{S}}=\{(x_{i},y_{i})|\hat{s}_{i}=1\}$, where $\hat{s}_{i}$ is the output of the slicing function on $i$-th sample, we define manifold compactness as:
@@
This metric is the average weighted degree of nodes of the induced subgraph, whose vertex set corresponds to the slice. The higher it is, the denser or more compact the subgraph is, implying a more coherent slice. Note that when applying this to evaluate multiple slice discovery algorithms, for convenience of comparison, we control the size of $\hat{\mathcal{S}}$ for those algorithms to be the same by taking the top $\alpha n$ data points sorted by the slicing function’s prediction probability. Here $n$ is the size of the dataset and $\alpha\in(0,1]$ is a fixed proportion. The operation of selecting data points with highest prediction probabilities is akin to calculating precision@$k$ in dcbench (Eyuboglu et al., 2022).
Figure 3: Percentage of increase of manifold compactness (“Comp.”) and decrease of variance (“Var.”) from coarse-grained slices to fine-grained ones in CelebA. For manifold compactness, there is always a positive increase from coarse-grained slices to fine-grained slices. For variance, there is not a monotonic trend. This could imply that manifold compactness is better at capturing semantic coherence than variance does.
@@
     $\hat{s}_{i}=1$ if $w_{i}>\alpha$-Quantile of $\{w_{i}\}_{i=1}^{n}$ else 00.
@@
Next, we try to demonstrate the validity and advantages of our proposed coherence metric. A common and representative metric of coherence directly calculated in Euclidean space is variance. Thus we measure variance and manifold compactness respectively on different semantically predefined slices of CelebA (Liu et al., 2015). We use the binary label $y$ to indicate whether the person has blond hair or not, and $a$ to indicate whether the person is male or not. The values of $y$ and $a$ can formulate slices of different granularity. In Figure 3, the most coarse-grained slice is the whole dataset (the darkest circle in the center), the most fine-grained slice is the combination of $y$ and $a$ (the lightest circles in the four corners), and slices of the middle granularity are formulated by either of $y$ and $a$. Figure 3 shows the percentage of the increase of manifold compactness and the decrease of variance with directed arrows from semantically coarse-grained slices to fine-grained ones. It is intuitive that these digits are supposed to be positive if these two metrics could properly measure semantic coherence. However, for variance, in some cases the value of the more coarse-grained slice is even smaller than the more fine-grained, marked in red arrows. For manifold compactness, there is always a positive increase from semantically coarse-grained slices to fine-grained slices. In this way, we demonstrate that manifold compactness is better at capturing semantic coherence than variance does. Still due to the space limit, we only provide results of CelebA here, and leave detailed values and results of other datasets in Section A.1.2, where we reach the same conclusion. Besides, in Table 1 of Section 5, we have also empirically shown that the rank order of the four methods according to precision metrics is generally the same as that of manifold compactness. Since the precision metrics are based on predefined slice labels with semantic meanings, it implies that our proposed coherence metric could capture semantic patterns well and is appropriate for evaluation of slice discovery algorithms even when predefined slice labels are absent.
4 Algorithm
@@
First, we approximate the data manifold via a graph. To facilitate the graph learning approach, we obtain the embeddings of the dataset via a pretrained feature extractor (Radford et al., 2021), i.e. $z_{i}=h_{fe}(x_{i})$, which follows previous works of error slice discovery (Eyuboglu et al., 2022; Wang et al., 2023b). Then we construct a kNN graph $G=(V,E,Q)$ based on the embeddings $\{z_{i}\}_{i=1}^{n}$, which is a widely adopted manifold learning approach (Zemel & Carreira-Perpiñán, 2004; Pedronette et al., 2018; Dann et al., 2022). In the graph $G$, the edge weight $q_{ij}=1$ if $z_{j}$ is among the $k$ nearest neighbors of $z_{i}$, or else $q_{ij}=0$.
@@
The above QP problem can be easily solved by classic optimization algorithms or powerful mathematical optimization solvers like Gurobi (Gurobi Optimization, 2021). After solving for the proper sample weights $\{w_{i}\}_{i=1}^{n}$, we select the top $\alpha n$ samples sorted by the weights as the error slice $\hat{\mathcal{S}}$. Note that in most previous algorithms’ workflow, they require the prediction probability as the input (Eyuboglu et al., 2022; Plumb et al., 2023; Wang et al., 2023b), thus only applicable to classification, while our algorithm takes the prediction loss as input, naturally more flexible and applicable to various tasks.

### v2

@@ Abstract
Despite the great performance of deep learning models in many areas, they still make mistakes and underperform on certain subsets of data, i.e. error slices. Given a trained model, it is important to identify its semantically coherent error slices that are easy to interpret, which is referred to as the error slice discovery problem. However, there is no proper metric of slice coherence without relying on extra information like predefined slice labels. Current evaluation of slice coherence requires access to predefined slices formulated by metadata like attributes or subclasses. Its validity heavily relies on the quality and abundance of metadata, where some possible patterns could be ignored. Besides, current algorithms cannot directly incorporate the constraint of coherence into their optimization objective due to absence of an explicit coherence metric, which could potentially hinder their effectiveness. In this paper, we propose manifold compactness, a coherence metric without reliance on extra information by incorporating the data geometry property into its design, and experiments on typical datasets empirically validate the rationality of the metric. Then we develop Manifold Compactness based error Slice Discovery (MCSD), a novel algorithm that directly treats risk and coherence as the optimization objective, and is flexible to be applied to models of various tasks. Extensive experiments on the benchmark and case studies on other typical datasets demonstrate the superiority of MCSD.
@@ 1 Introduction
In recent years, with the enhancement of computational power, neural networks have achieved significant progress in numerous tasks (Achiam et al. 2023; Liu et al. 2023a; Tian, Ye, and Doermann 2025). Despite their impressive overall performance, they are far from perfect, and still suffer from performance degradation on some subpopulations (Yang et al. 2023). This substantially hinders their application in risk-sensitive scenarios like medical imaging (Yang et al. 2024), autonomous driving (Chen et al. 2024), etc., where model mistakes may result in catastrophic consequences. Therefore, to avoid the misuse of models, it is a fundamental problem to identify subsets (or slices) where a given model tends to underperform. Moreover, we would like to find coherent interpretable semantic patterns in the underperforming slices. For example, a facial recognition model may underperform in certain demographic groups like elderly females. An autonomous driving system may fail in the face of steep road conditions. Identifying such coherent patterns could help us understand model failures, and we could employ straightforward solutions for improvement like collecting new data (Liu et al. 2023b) or upweighting samples in error slices (Liu et al. 2021a).
@@
Previously, works of error slice discovery (d’Eon et al. 2022; Wang et al. 2023) aim for this goal. Despite the emphasis on coherence in error slice discovery, there is no proper metric to assess the coherence of a given slice without additional information like predefined slice labels. On one hand, this impairs the efficacy of the evaluation paradigm of error slice discovery. In the current benchmark (Eyuboglu et al. 2022), with the help of metadata like attributes or subclasses, it predefines slices that are already semantically coherent, and they depict the coherence of a slice discovered by a specific algorithm via the matching degrees between it and the predefined underperforming slices, so as to evaluate the effectiveness of the algorithm. Such practice heavily relies on not only the availability but also the quality of metadata, whose annotations are usually expensive, and may overlook model failure patterns not captured by existing metadata. On the other hand, due to the absence of an explicit coherence metric, current algorithms can only indirectly incorporate the constraint of coherence into their design, e.g. via clustering (Eyuboglu et al. 2022; Wang et al. 2023; Plumb et al. 2023), without treating it as a direct optimization objective. This could potentially impede the development of more effective error slice discovery algorithms.
@@
Figure 1: Illustration of MCSD. Blue points are correctly classified by the given trained model, while red ones are wrongly classified. The trained model achieves a good overall accuracy, but exhibit a high error in a certain slice.
@@
In this paper, inspired by the data geometry property that high dimensional data tends to lie on a low-dimensional manifold (Belkin and Niyogi 2003; Roweis and Saul 2000; Tenenbaum, Silva, and Langford 2000), we incorporate this property to propose manifold compactness as the metric of coherence given a slice, which does not require additional information. We illustrate the validity of the metric by showing that it captures semantic patterns better than depicting coherence via metrics directly calculated in Euclidean space, and is empirically consistent with current evaluation metrics that require predefined slice labels. Then we propose a novel and flexible algorithm named Manifold Compactness based error Slice Discovery (MCSD) that jointly optimizes the average risk and manifold compactness to identify the error slice. Thus both the risk and coherence, i.e. the desired properties of error slices are explicitly treated as the optimization objective. We illustrate our algorithm in Figure 1. Besides, our algorithm can be directly applied to trained models of different tasks while most error slice discovery methods are restricted to classification only. We provide theoretical analyses of the algorithm. We conduct experiments on dcbench (Eyuboglu et al. 2022) to demonstrate our algorithm’s superiority compared with existing ones. We also provide several case studies on different types of datasets and tasks to showcase the effectiveness and flexibility of our algorithm. Our contributions are summarized below:
@@ 2 Problem
Due to space limit, we leave the section of related works in Appendix B. Unless stated otherwise, for random variables, we use uppercase letters, in contrast to a concrete dataset where we use lowercase letters. Consider classic supervised learning. The input variable is denoted as $X\in\mathcal{X}$ and the outcome is denoted as $Y\in\mathcal{Y}$, whose joint distribution is $P(X,Y)$. There exist multiple slices, where $j$-th slice can be represented as a slice label variable $S^{(j)}\in\{0,1\}$. For classic supervised learning, the goal is to learn a model $f_{\theta}:\mathcal{X}\mapsto\mathcal{Y}$ with parameter $\theta$. Denote $\ell:\mathcal{Y}\times\mathcal{Y}\mapsto[0,+\infty]$ as the loss function. Current machine learning algorithms are capable of learning models with a satisfying overall performance, which can be demonstrated via a low risk $\mathbb{E}_{P}[\ell(f_{\theta}(X),Y)]$ over the whole population. However, performance degradation could still occur in a certain subpopulation or slice. Here we introduce error slice discovery:
@@
The reason why we require an extra validation dataset to implement error slice discovery is that for deep learning models, training data is usually fitted well enough or even nearly perfect. Thus model mistakes on training data carry much less information on models’ generalization ability. This is common practice in previous works (d’Eon et al. 2022; Eyuboglu et al. 2022; Wang et al. 2023). Without ambiguity, we omit the superscript or subscript of “va” for $n,x_{i},y_{i}$ for convenience in the next two sections.
@@ 3 Metric
Due to the absence of a proper metric for coherence that is independent of additional information, the current benchmark (Eyuboglu et al. 2022) provides numerous datasets, trained models, and their predefined underperforming slice labels. They employ precision@$k$, i.e. the proportion of the top $k$ elements in the discovered slice belonging to the predefined ground-truth error slice as the metric of slice coherence to evaluate error slice discovery algorithms. Although such practice is reasonable to some extent, its effectiveness of evaluation strongly relies on the quality of metadata that composes the underperforming slice labels, which might be not even available under many circumstances.
@@
Figure 2: Category “Blond Hair” of CelebA. Visualization of t-SNE and UMAP (manifold-based dimension reduction) shows much clearer clustering structures than that of PCA (mainly preserving Euclidean distances between data points). Thus it could be better to measure coherence in the metric space of a manifold than using metrics directly calculated in Euclidean space.
To eliminate the requirement of predefined slices, we try to propose a new metric of coherence. It is commonly acknowledged that high-dimensional data usually lies on a low-dimensional manifold (Belkin and Niyogi 2003; Roweis and Saul 2000; Tenenbaum, Silva, and Langford 2000). In this case, while direct usage of Euclidean distance cannot properly capture the dissimilarity between data points, the geodesic distance in the metric space of the manifold can. For preliminary justification, here we provide visualization analyses based on different types of dimension-reduction techniques. Among these techniques, PCA mainly preserves pairwise Euclidean distances between data points while t-SNE and UMAP are both manifold learning techniques. In Figure 2, blue dots are correctly classified by the trained model and red dots are wrongly classified. We can see that the visualization of t-SNE and UMAP shows much clearer clustering structures than that of PCA, either having a larger number of clusters or exhibiting larger margins between clusters. This indicates that it could be better to measure coherence in the metric space of a manifold than in the original Euclidean space. Due to space limit, we only present results of the widely adopted facial dataset CelebA (Liu et al. 2015) here, leaving results of other datasets in Section A.1, where the same conclusion holds.
Therefore, we attempt to define a metric of coherence inside the discovered slice via compactness in the data manifold. In practice, a manifold can be treated as a graph $G$ (Melas-Kyriazi 2020), and we can apply graph learning methods like k-nearest neighbor (kNN) to approximate it (Dann et al. 2022). Given an identified slice $\hat{\mathcal{S}}=\{(x_{i},y_{i})|\hat{s}_{i}=1\}$, where $\hat{s}_{i}$ is the output of the slicing function on $i$-th sample, we define manifold compactness as:
@@
This metric is the average weighted degree of nodes of the induced subgraph, whose vertex set corresponds to the slice. The higher it is, the denser or more compact the subgraph is, implying a more coherent slice. Note that when applying this to evaluate multiple slice discovery algorithms, for convenience of comparison, we control the size of $\hat{\mathcal{S}}$ for those algorithms to be the same by taking the top $\alpha n$ data points sorted by the slicing function’s prediction probability. Here $n$ is the size of the dataset and $\alpha\in(0,1]$ is a fixed proportion. The operation of selecting data points with highest prediction probabilities is akin to calculating precision@$k$ in dcbench (Eyuboglu et al. 2022).
Figure 3: Percentage of increase of manifold compactness (“Comp.”) and decrease of variance (“Var.”) from coarse-grained slices to fine-grained ones in CelebA. For manifold compactness, there is always a positive increase from coarse-grained slices to fine-grained slices. However, in some cases, variance fails to decrease from coarse-grained slices to fine-grained slices as expected, which are marked in red arrows. This could imply that manifold compactness is better at capturing semantic coherence than variance does.
Next, we try to demonstrate the validity and advantages of our proposed coherence metric. A common and representative metric of coherence directly calculated in Euclidean space is variance. Thus we measure variance and manifold compactness respectively on different semantically predefined slices of CelebA (Liu et al. 2015). We use the binary label $y$ to indicate whether the person has blond hair or not, and $a$ to indicate whether the person is male or not. The values of $y$ and $a$ can formulate slices of different granularity. In Figure 3, the most coarse-grained slice is the whole dataset (the darkest circle in the center), the most fine-grained slice is the combination of $y$ and $a$ (the lightest circles in the four corners), and slices of the middle granularity are formulated by either of $y$ and $a$. Figure 3 shows the percentage of the increase of manifold compactness and the decrease of variance with directed arrows from semantically coarse-grained slices to fine-grained ones. It is intuitive that these digits are supposed to be positive if these two metrics could properly measure semantic coherence. However, for variance, in some cases the value of the more coarse-grained slice is even smaller than the more fine-grained, marked in red arrows. For manifold compactness, there is always a positive increase from semantically coarse-grained slices to fine-grained slices. In this way, we demonstrate that manifold compactness is better at capturing semantic coherence than variance does. Still due to the space limit, we only provide results of CelebA here, and leave detailed values and results of other datasets, along with comparisons against other metrics directly calculated in Euclidean space like Median Absolute Deviation (MedianAD) in Section A.1, where we reach the same conclusion. Besides, in Table 1 of Section 5, we have also empirically shown that the rank order of the four methods according to precision metrics is generally the same as that of manifold compactness. Since the precision metrics are based on predefined slice labels with semantic meanings, it implies that our proposed coherence metric could capture semantic patterns well and is appropriate for evaluation of slice discovery algorithms even when predefined slice labels are absent.
4 Algorithm
@@
  $\hat{s}_{i}=1$ if $w_{i}>\alpha$-Quantile of $\{w_{i}\}_{i=1}^{n}$ else 0.
@@
First, we approximate the data manifold via a graph. To facilitate the graph learning approach, we obtain the embeddings of the dataset via a pretrained feature extractor (Radford et al. 2021), i.e. $z_{i}=h_{fe}(x_{i})$, which follows previous works of error slice discovery (Eyuboglu et al. 2022; Wang et al. 2023). Then we construct a kNN graph $G=(V,E,Q)$ based on the embeddings $\{z_{i}\}_{i=1}^{n}$, which is a widely adopted manifold learning approach (Zemel and Carreira-Perpiñán 2004; Pedronette, Gonçalves, and Guilherme 2018; Dann et al. 2022). In the graph $G$, the edge weight $q_{ij}=1$ if $z_{j}$ is among the $k$ nearest neighbors of $z_{i}$, or else $q_{ij}=0$.
@@
The above QP problem can be solved by classic optimization algorithms or powerful mathematical optimization solvers like Gurobi (Gurobi Optimization 2021). After solving for the proper sample weights $\{w_{i}\}_{i=1}^{n}$, we select the top $\alpha n$ samples sorted by the weights as the error slice $\hat{\mathcal{S}}$. Note that in most previous algorithms’ workflow, they require prediction probabilities as input (Eyuboglu et al. 2022; Plumb et al. 2023; Wang et al. 2023), thus only applicable to classification, while our algorithm takes prediction loss as input, naturally more flexible and applicable to various tasks.